2

Consider an economy that produces only three types of fruit: apples bananas, and oranges. In the base year (a few years ago), the production and price data are listed in the tables to the right. Base Year Quantity Fruit Apples Bananas 5,000 bunches Oranges Price 3,000 bags In the base year, nominal GUP was s In the current year, nominal GDP is $ (Round both answers to the nearest whole number.) $2 per bag $3 per bunch $5 per bag 6,000 bags The percentage increase in nominal GDP since the base year is 1 % (Enter your response as a percentage rounded to one decimal place.) Current Year Quantity Fruit Apples Bananas 13,000 bunches Oranges Price 4,000 bags In the base year, real GDP was $ In the current year, real GDP is (Round both answers to the nearest whole number.) $3 per bag $2 per bunch $7 per bag 24,000 bags The percentage increase in real GDP since the base year is | |%. (Enter your response as a percentage rounded to one decimal place.) The GDP deflator for the base year is The GDP deflator for the current year is (Round both answers to one decimal place.) The inflation rate (which equals the percentage increase in the GDP deflator) since the base year is 1 %. Enter your response as a percentage rounded to one decimal place.) The percentage increase in nominal GDP in this economy since the base year is due primarily to

3
Answer #1

2.

Base Year

Fruit

Quantity

Price

Apples

3000

2

Bananas

5000

3

Oranges

6000

5

Nominal GDP in Base Year

51000

Current Year

Fruit

Quantity

Price

Apples

4000

2

Bananas

13000

3

Oranges

24000

7

Nominal GDP in Current year

215000

The percentage increase in Nominal GDP is ((215000-51000)/51000)*100 = 321.6 percent

In the base year, real GDP was 51000

Current Year

Fruit

Quantity

Price

Apples

4000

2

Bananas

13000

3

Oranges

24000

5

Nominal GDP

167000

In the base year, real GDP was 167000.

The percentage increase in Real GDP is ((167000-51000)/51000)*100 = 227.5 percent

The GDP deflator for the base year is 100

The GDP deflator for the current year is Nominal GDP/Real GDP = (215000/167000)*100 =128.7

The inflation rate is 128.7-100 = 28.7 percent

Is primarily due to increase in price of Oranges

3.

Year

CPI

Bond Value

Nominal Interest rate

2015

200

500

2016

202

525

=(202-200))/200)*100=5%

Actual inflation rate = ((202-100)/100)*100 = 102 percent

Real interest rate = Nominal – Inflation = 5%-2% =3%

Expected inflation rate = 1%

Expected real interest rate = 5%-1%=4%